OpenStudy (anonymous):

2x(3x - 4) = 3x(2x + 1) - 11 Does this have a positive or negative outcome

OpenStudy (anonymous):

First solve it: Open the brackets. 6x^2 -8x=6x^2-+3x-11 Move everything to left side: 6x^2-6x^2-8x-3x+11=0 Solve -11x+11=0 x=1 so positive

OpenStudy (mayankdevnani):

ya!! positive...

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

No problem!